# Mathematical Language Models: A Survey

### 14 Dec 2023 v2

@@ Abstract
[…]
1 Introduction
[…]
@@ 1 Introduction
Mathematics stands as a foundational skill integral to human intelligence, wielding significance across diverse fields such as natural sciences, engineering, medicine, finance, computer science, and social sciences. Within the field of natural language processing (NLP), the development of computer models aimed at autonomously resolving mathematical word problems has captivated researchers since as early as 1963 [feigenbaum1963computers, bobrow1964natural, briars1984integrated, fletcher1985understanding]. We believe that addressing this problem is a potential way toward illuminating pathways for general reasoning mechanisms, consequently advancing the pursuit of general artificial intelligence (AI).
@@
One main traditional solution to math word problems is statistical learning-based methodologies [hosseini-etal-2014-learning, kushman-etal-2014-learning, zhou2015learn, mitra2016learning]. Notably, machine learning techniques [kushman-etal-2014-learning, roy2015reasoning, roy2015solving] alongside semantic parsing methods [shi2015automatically, koncel2015parsing] have been employed to address this challenge, showcasing promising outcomes on certain datasets. The evolution of deep learning has prompted considerable interest in crafting neural networks capable of resolving mathematical problems [wang2017deep, couperus2023large]. Notably, recent years have witnessed significant advancements in the domain of mathematical AI [wei2022chain, matzakos2023learning], propelled by the emergence of powerful language models (LMs) [kenton2019bert, touvron2023llama]. These language models, comprising pre-trained language models (PLMs) and large-scale language models (LLMs), have assumed a central role in reshaping the landscape of mathematical exploration and practical applications. The paramount focus of this comprehensive survey lies in assessing the impact of these models on the field of mathematics. The survey endeavors to provide a thorough overview of extant research and its consequential implications.
@@
PLMs such as BERT [devlinBERTPretrainingDeep2019], RoBERTa [liuRoBERTaRobustlyOptimized2019], BART [lewis2020bart], GPT-1 [radford_gpt-1_2018] and GPT-2 [radford_gpt-2_2019] undergo pre-training on extensive textual corpora to assimilate worldly knowledge. To enhance mathematical performance, certain endeavors focus on either pre-training or fine-tuning PLMs using mathematical datasets [clarkRegentsScienceExams2021, gevaInjectingNumericalReasoning2020, fengInjectingNumericalReasoning2022]. For instance, methodologies like GenBERT [gevaInjectingNumericalReasoning2020], NF-NSM [fengInjectingNumericalReasoning2022], MathBERT[pengMathBERTPretrainedModel2021] and LISA [jiang2021lisa] incorporate numerical data or mathematical formulas into PLMs to augment their capabilities. Moreover, specialized modules or tailored loss functions are devised, leveraging existing PLMs to learn mathematical reasoning and operations [zhu2021tat, zhao2022multihiertt, jieLearningReasonDeductively2022, liSeekingPatternsNot2022].
@@
The recent advent of LLMs, exemplified by OpenAI’s GPT-4 [openai2023gpt4], has catalyzed an unforeseen surge in innovation, underscoring the multifaceted potential of AI within the domain of mathematics. These models ([touvron2023llama, touvron2023llama2]) have demonstrated remarkable success across diverse Natural Language Processing (NLP) tasks, leveraging in-context learning [min2022metaicl, brown2020language, chen2022meta] and instruction learning [yang2023gpt, liu_goat_2023]. Recent studies by Wang et al. [wang2022self] indicate that LLMs featuring over 100 billion parameters (e.g., GPT-3 [brown2020language] with 175 billion, PaLM [chowdhery2022palm] with 540 billion) exhibit the capability to address intricate tasks by employing a chain-of-thought (CoT) mechanism [wei2022chain] when furnished with a limited set of reasoning examples as demonstrations. Advancements in CoT frameworks [chen2022program, lu2022dynamic, zhang2023automatic, fu2022complexity] have been tailored to enhance mathematical performance, incorporating tools and programs [gao2023pal, drori2022neural, he2023solving]. An exemplar in this domain is LLEMMA [azerbayev8llemma], an openly accessible LLM specifically designed for mathematical tasks. LLEMMA showcases the capacity to generate self-contained textual solutions for mathematical problems and formulate formal proofs.
@@
We mainly review the most related surveys about mathematical reasoning and LLMs. Lu et al. [lu-etal-2023-survey] delineate the landscape of deep learning applications specifically pertaining to mathematical reasoning. Qiao et al. [qiao-etal-2023-reasoning] predominantly delve into reasoning mechanisms centered around language models, encompassing arithmetic, commonsense, logical, symbolic, and multi-modal reasoning. Chu et al. [chuCoTReasoningSurvey2023] offer an exploration of studies focused on chain-of-thought (CoT) reasoning. Additionally, Qiu el al. [qiu2020pre] and Naveed et al. [naveed2023comprehensive] contribute comprehensive reviews elucidating the realm of existing PLMs and LLMs, respectively.
@@ 2 Mathematical Tasks
Representation [ DigitRNN  [spithourakis2018numeracy], DigitCNN  [wallace2019nlp], GenBERT  [gevaInjectingNumericalReasoning2020], NumBERT  [zhangLanguageEmbeddingsCapture2020] , leaf, text width=44.6em ] ] [ Arithmetic
Calculation [ ScratchpadGPT  [nye_show_2021], Goat  [liu_goat_2023], MathGLM  [yang2023gpt], , leaf, text width=44.6em ] ] ] [ Mathematical
@@
Solving [ MathPrompter  [imani2023mathprompter], MetaMath  [yu2023metamath], WizardMath  [luo2023wizardmath], MathAttack  [zhou2023mathattack], LLEMMA  [azerbayev8llemma] , leaf, text width=44.6em ] ] [ Math Question
Answering [ GEOS  [seo-etal-2015-solving], DROP  [dua2019drop], Mathematics  [saxton2018analysing], Lila  [mishra2022lila] , leaf, text width=44.6em ] ] [ Theorem
Proving [ DeepMath  [irving2016deepmath], ASTactic  [yang2019learning], NaturalProofs  [welleck2021naturalproofs], INT  [wu2020int] , leaf, text width=44.6em ] ] ] ]
Figure 1: Taxonomy of mathematical tasks.
2.1 Mathematical Calculation
@@
2.1.1 Arithmetic Representation
In the early stages of the research, numerical values were either omitted or oversimplified, treated as ordinary text, or categorized as “unknown" (UNK). This approach, however, proved inadequate for tasks. For instance, BERT performs five times worse when the answer is a number rather than a textual span over the DROP question-answering benchmark [dua2019drop].
Recent literature proposes various methods for numerical representation. Geva et al. [gevaInjectingNumericalReasoning2020] introduce GenBERT, which tokenizes numbers at the digit level and undergoes fine-tuning with arithmetic word problems and simple arithmetic tasks. Zhang et al. [zhang2020language] experiment with a retrained BERT model by converting numbers into scientific notation (e.g., 314.1 as 3141[EXP]2). Spithourakis and Riedel [spithourakis2018numeracy], along with Wallace et al. [wallace2019nlp], explore the integration of digit embeddings into a singular embedding that represents the entire number. Berg-Kirkpatrick and Spokoyny [berg-kirkpatrickEmpiricalInvestigationContextualized2020] propose a method with digit-RNN and exponent embeddings. This method specifically emphasizes the exponent while disregarding the mantissa. Goat [liu_goat_2023] introduces Consistent Tokenization, enhancing the relationships between similar numerical values.
2.1.2 Arithmetic Calculation
There has been significant research into the arithmetic capabilities of LLMs. Nogueira et al. [nogueira2021investigating] and Wang et al. [wang2021exploring] assessed addition and subtraction tasks. Muffo et al. [muffo_evaluating_nodate] evaluated two-digit multiplication. Yuan et al. [yuan2023well] evaluated the arithmetic operation capabilities of different models, including GPT-4 [openai2023gpt4], Galactica [taylor_galactica_2022], and LLaMA [touvron2023llama].
Traditionally, it was presumed that LLMs could not accurately perform complex arithmetic, especially multiplication involving more than eight digits. However, recent approaches challenge this assumption. Zhou et al. [zhouTeachingAlgorithmicReasoning2022] apply specialized prompt engineering to enhance addition capabilities but note limitations in multiplication beyond seven digits. Jelassi et al. [jelassiLengthGeneralizationArithmetic2023] investigate length generalization in basic arithmetic tasks using techniques like relative position embeddings and training set priming. ScratchpadGPT [nye_show_2021] demonstrates the effectiveness of pre-generating a Chain of Thought (CoT) before producing an answer in 8-digit addition tasks. Goat [liu_goat_2023] utilizes supervised instruction for fine-tuning elementary arithmetic operations with large integers. MathGLM [yang2023gpt] excels in intricate arithmetic tasks by pre-training on datasets that decompose complex arithmetic expressions into simpler steps. Through this process, it gradually generates answers and learns the rules of computation.
2.2 Mathematical Reasoning
@@
2.2.1 Math Problem Solving
@@
Math Word Problems (MWPs) refer to problems that present mathematical concepts and calculations in the form of written descriptions. Such questions usually include one or more story situations describing a series of events or conditions from which the solver needs to extract relevant mathematical information and apply appropriate mathematical principles to solve the problem. Over the years, using various efficient and intelligent algorithms to solve MWPs has also been a research focus in the fields of computer science and artificial intelligence [bobrow1964natural].
@@
Recent research has highlighted the growing capabilities of LLMs in the field of mathematical word problem solving, emphasizing the trend toward more nuanced and sophisticated AI-driven mathematical analysis. MathPrompter [imani2023mathprompter] uses LLM called GPT3 DaVinci to solve MWPs with excellent results, demonstrating the potential of LLMs to not only explain but also generate complex mathematical reasoning, reflecting a human-like understanding of complex problem sets.
@@
Yuan et al. [yuan2023scaling] investigated the interaction between various factors such as pre-training loss, amount of supervised data, and augmented data to improve the mathematical inference performance of LLMs. They found that pre-training loss is a more reliable performance metric and empirically found a log-linear relation between data amount and model performance, but it is worth mentioning that this improvement diminishes with better pre-training models. At the same time, they proposed Rejection sampling Fine-Tuning (RFT) to enhance performance by using supervised models to generate and collect correct reasoning paths, thereby achieving better inference generalization and more significant improvements. Meanwhile, the proficiency with which LLMs perform basic arithmetic tasks is critical to establishing the baseline capabilities of these models [yuan2023well]. This assessment is integral to understanding the limitations and potential of LLMs in future work when faced with mathematical challenges that require precise and logical calculations.
@@
MetaMath [yu2023metamath] further extends the practicality of the LLMs, proposing a paradigm whereby the LLMs generate their mathematical problems, thereby creating a self-sustaining learning environment that encourages continuous improvement of the model’s problem-solving acuity. Similarly, WizardMath [luo2023wizardmath] explores new ways to enhance LLMs’ mathematical reasoning capabilities by reinforcing evolutionary instructions, indicating an important step towards autonomous self-improvement of artificial intelligence models. Instead, MathAttack [zhou2023mathattack] introduces a critical perspective by exploring the sensitivity of LLMs to specialized adversarial input designed to test their mathematical problem-solving abilities. Such investigations are crucial to developing more resilient and reliable LLMs that can withstand and adapt to a variety of problem-solving situations. LLEMMA [azerbayev8llemma], an open language model dedicated to mathematics, outlines ongoing efforts to make LLMs accessible and adaptable to a wide range of mathematical inquiries, from basic problem solving to advanced theorem proving.
@@
Sachan et al. [saxton2018analysing] introduce a framework for evaluating neural architectures through a suite of math problems, highlighting the importance of algebraic generalization. This is complemented by GEOS [seo-etal-2015-solving], an innovative system that solves SAT geometry problems by integrating text understanding with diagram interpretation. The introduction of datasets and systems like Inter-GPS [lu2021inter], IconQA [lu2021iconqa], and PGDP5K [Hao2022PGDP5KAD, Zhang2022, Zhang2023PGPS] further advances the field, offering new benchmarks for abstract diagram understanding and visual language reasoning. Additionally, Scharpf et al. [scharpf2022mining] indicate that unsupervised methods for formula labeling contribute to the automated understanding of mathematical documents. Collectively, these works demonstrate the evolving landscape of MQA, where the synergy between natural language processing, computer vision, and symbolic reasoning opens new avenues for educational technology and AI research.
2.2.2 Theorem Proving

### 23 Feb 2024 v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Natural language generationccs: Computing methodologies Scene understandingccs: Computing methodologies Cognitive roboticsccs: Computing methodologies Cognitive scienceccs: Computing methodologies Intelligent agents
1. Introduction
@@
Mathematics stands as a foundational skill integral to human intelligence, wielding significance across diverse fields such as natural sciences, engineering, medicine, finance, computer science, and social sciences. Within the field of natural language processing (NLP), the development of computer models aimed at autonomously resolving mathematical word problems has captivated researchers since as early as 1963 (Feigenbaum et al., 1963; Bobrow et al., 1964; Briars and Larkin, 1984; Fletcher, 1985). We believe that addressing this problem is a potential way toward illuminating pathways for general reasoning mechanisms, consequently advancing the pursuit of general artificial intelligence (AI).
@@
One main traditional solution to math word problems is statistical learning-based methodologies (Hosseini et al., 2014; Kushman et al., 2014; Zhou et al., 2015; Mitra and Baral, 2016). Notably, machine learning techniques (Kushman et al., 2014; Roy et al., 2015; Roy and Roth, 2015) alongside semantic parsing methods (Shi et al., 2015; Koncel-Kedziorski et al., 2015) have been employed to address this challenge, showcasing promising outcomes on certain datasets. The evolution of deep learning has prompted considerable interest in crafting neural networks capable of resolving mathematical problems (Wang et al., 2017; Couperus, 2023). Notably, recent years have witnessed significant advancements in the domain of mathematical AI (Wei et al., 2022; Matzakos et al., 2023), propelled by the emergence of powerful language models (LMs) (Kenton and Toutanova, 2019; Touvron et al., 2023a). These language models, comprising pre-trained language models (PLMs) and large-scale language models (LLMs), have assumed a central role in reshaping the landscape of mathematical exploration and practical applications. The paramount focus of this comprehensive survey lies in assessing the impact of these models on the field of mathematics. The survey endeavors to provide a thorough overview of extant research and its consequential implications.
@@
PLMs such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020), GPT-1 (Radford et al., 2018) and GPT-2 (Radford et al., 2019) undergo pre-training on extensive textual corpora to assimilate worldly knowledge. To enhance mathematical performance, certain endeavors focus on either pre-training or fine-tuning PLMs using mathematical datasets (Clark et al., 2021; Geva et al., 2020; Feng et al., 2022). For instance, methodologies like GenBERT (Geva et al., 2020), NF-NSM (Feng et al., 2022), MathBERT(Peng et al., 2021) and LISA (Jiang et al., 2021) incorporate numerical data or mathematical formulas into PLMs to augment their capabilities. Moreover, specialized modules or tailored loss functions are devised, leveraging existing PLMs to learn mathematical reasoning and operations (Zhu et al., 2021; Zhao et al., 2022; Jie et al., 2022; Li et al., 2022).
@@
The recent advent of LLMs, exemplified by OpenAI’s GPT-4 (OpenAI, 2023), has catalyzed an unforeseen surge in innovation, underscoring the multifaceted potential of AI within the domain of mathematics. These models ((Touvron et al., 2023a, b)) have demonstrated remarkable success across diverse Natural Language Processing (NLP) tasks, leveraging in-context learning (Min et al., 2022; Brown et al., 2020; Chen et al., 2022b) and instruction learning (Yang et al., 2023; Liu and Low, 2023). Recent studies by Wang et al. (Wang et al., 2023b) indicate that LLMs featuring over 100 billion parameters (e.g., GPT-3 (Brown et al., 2020) with 175 billion, PaLM (Chowdhery et al., 2022) with 540 billion) exhibit the capability to address intricate tasks by employing a chain-of-thought (CoT) mechanism (Wei et al., 2022) when furnished with a limited set of reasoning examples as demonstrations. Advancements in CoT frameworks (Chen et al., 2022a; Lu et al., 2023b; Zhang et al., 2023c; Fu et al., 2023) have been tailored to enhance mathematical performance, incorporating tools and programs (Gao et al., 2023; Drori et al., 2022; He-Yueya et al., 2023). An exemplar in this domain is LLEMMA (Azerbayev et al., [n. d.]), an openly accessible LLM specifically designed for mathematical tasks. LLEMMA showcases the capacity to generate self-contained textual solutions for mathematical problems and formulate formal proofs.
@@
We mainly review the most related surveys about mathematical reasoning and LMs (Min et al., 2023). Lu et al. (Lu et al., 2023c) delineate the landscape of deep learning applications specifically pertaining to mathematical reasoning. Qiao et al. (Qiao et al., 2023) predominantly delve into reasoning mechanisms centered around language models, encompassing arithmetic, commonsense, logical, symbolic, and multi-modal reasoning. Chu et al. (Chu et al., 2023) offer an exploration of studies focused on chain-of-thought (CoT) reasoning. Additionally, Qiu el al. (Qiu et al., 2020) and Zhao et al. (Zhao et al., 2023b) contribute comprehensive reviews elucidating the realm of existing PLMs and LLMs, respectively.
@@ 2. Mathematical Tasks
Representation [ DigitRNN  (Spithourakis and Riedel, 2018), DigitCNN  (Wallace et al., 2019), GenBERT  (Geva et al., 2020), NumBERT  (Zhang et al., 2020a) , leaf, text width=44.6em ] ] [ Arithmetic
Calculation [ ScratchpadGPT  (Nye et al., 2021), Goat  (Liu and Low, 2023), MathGLM  (Yang et al., 2023), , leaf, text width=44.6em ] ] ] [ Mathematical
@@
Solving [ MathPrompter  (Imani et al., 2023), MetaMath  (Yu et al., 2023b), WizardMath  (Luo et al., 2023), MathAttack  (Zhou et al., 2023c), LLEMMA  (Azerbayev et al., [n. d.]) , leaf, text width=44.6em ] ] [ Math Question
Answering [ GEOS  (Seo et al., 2015), DROP  (Dua et al., 2019), Mathematics  (Saxton et al., 2019), Lila  (Mishra et al., 2022a) , leaf, text width=44.6em ] ] [ Theorem
Proving [ DeepMath  (Irving et al., 2016), ASTactic  (Yang and Deng, 2019), NaturalProofs  (Welleck et al., 2021), INT  (Wu et al., 2021) , leaf, text width=44.6em ] ] ] ]
Figure 1. Taxonomy of mathematical tasks.
2.1. Mathematical Calculation
@@
2.1.1. Arithmetic Representation
In the early stages of the research, numerical values were either omitted or oversimplified, treated as ordinary text, or categorized as “unknown” (UNK). This approach, however, proved inadequate for tasks. For instance, BERT performs five times worse when the answer is a number rather than a textual span over the DROP question-answering benchmark (Dua et al., 2019).
Recent literature proposes various methods for numerical representation. Geva et al. (Geva et al., 2020) introduce GenBERT, which tokenizes numbers at the digit level and undergoes fine-tuning with arithmetic word problems and simple arithmetic tasks. Zhang et al. (Zhang et al., 2020b) experiment with a retrained BERT model by converting numbers into scientific notation (e.g., 314.1 as 3141[EXP]2). Spithourakis and Riedel (Spithourakis and Riedel, 2018), along with Wallace et al. (Wallace et al., 2019), explore the integration of digit embeddings into a singular embedding that represents the entire number. Berg-Kirkpatrick and Spokoyny (Berg-Kirkpatrick and Spokoyny, 2020) propose a method with digit-RNN and exponent embeddings. This method specifically emphasizes the exponent while disregarding the mantissa. Goat (Liu and Low, 2023) introduces Consistent Tokenization, enhancing the relationships between similar numerical values.
2.1.2. Arithmetic Calculation
There has been significant research into the arithmetic capabilities of LLMs. Nogueira et al. (Nogueira et al., 2021) and Wang et al. (Wang et al., 2021) assessed addition and subtraction tasks. Muffo et al. (Muffo et al., [n. d.]) evaluated two-digit multiplication. Yuan et al. (Yuan et al., 2023b) evaluated the arithmetic operation capabilities of different models, including GPT-4 (OpenAI, 2023), Galactica (Taylor et al., 2022), and LLaMA (Touvron et al., 2023a).
Traditionally, it was presumed that LLMs could not accurately perform complex arithmetic, especially multiplication involving more than eight digits. However, recent approaches challenge this assumption. Zhou et al. (Zhou et al., 2022) apply specialized prompt engineering to enhance addition capabilities but note limitations in multiplication beyond seven digits. Jelassi et al. (Jelassi et al., 2023) investigate length generalization in basic arithmetic tasks using techniques like relative position embeddings and training set priming. ScratchpadGPT (Nye et al., 2021) demonstrates the effectiveness of pre-generating a Chain of Thought (CoT) before producing an answer in 8-digit addition tasks. Goat (Liu and Low, 2023) utilizes supervised instruction for fine-tuning elementary arithmetic operations with large integers. MathGLM (Yang et al., 2023) excels in intricate arithmetic tasks by pre-training on datasets that decompose complex arithmetic expressions into simpler steps. Through this process, it gradually generates answers and learns the rules of computation.
2.2. Mathematical Reasoning
@@
2.2.1. Math Problem Solving
@@
Math Word Problems (MWPs) refer to problems that present mathematical concepts and calculations in the form of written descriptions. Such questions usually include one or more story situations describing a series of events or conditions from which the solver needs to extract relevant mathematical information and apply appropriate mathematical principles to solve the problem. Over the years, using various efficient and intelligent algorithms to solve MWPs has also been a research focus in the fields of computer science and artificial intelligence (Bobrow et al., 1964).
@@
Recent research has highlighted the growing capabilities of LLMs in the field of mathematical word problem solving, emphasizing the trend toward more nuanced and sophisticated AI-driven mathematical analysis. MathPrompter (Imani et al., 2023) uses LLM called GPT3 DaVinci to solve MWPs with excellent results, demonstrating the potential of LLMs to not only explain but also generate complex mathematical reasoning, reflecting a human-like understanding of complex problem sets.
@@
Yuan et al. (Yuan et al., 2023a) investigated the interaction between various factors such as pre-training loss, amount of supervised data, and augmented data to improve the mathematical inference performance of LLMs. They found that pre-training loss is a more reliable performance metric and empirically found a log-linear relation between data amount and model performance, but it is worth mentioning that this improvement diminishes with better pre-training models. At the same time, they proposed Rejection sampling Fine-Tuning (RFT) to enhance performance by using supervised models to generate and collect correct reasoning paths, thereby achieving better inference generalization and more significant improvements. Meanwhile, the proficiency with which LLMs perform basic arithmetic tasks is critical to establishing the baseline capabilities of these models (Yuan et al., 2023b). This assessment is integral to understanding the limitations and potential of LLMs in future work when faced with mathematical challenges that require precise and logical calculations.
@@
MetaMath (Yu et al., 2023b) further extends the practicality of the LLMs, proposing a paradigm whereby the LLMs generate their mathematical problems, thereby creating a self-sustaining learning environment that encourages continuous improvement of the model’s problem-solving acuity. Similarly, WizardMath (Luo et al., 2023) explores new ways to enhance LLMs’ mathematical reasoning capabilities by reinforcing evolutionary instructions, indicating an important step towards autonomous self-improvement of artificial intelligence models. Instead, MathAttack (Zhou et al., 2023c) introduces a critical perspective by exploring the sensitivity of LLMs to specialized adversarial input designed to test their mathematical problem-solving abilities. Such investigations are crucial to developing more resilient and reliable LLMs that can withstand and adapt to a variety of problem-solving situations. LLEMMA (Azerbayev et al., [n. d.]), an open language model dedicated to mathematics, outlines ongoing efforts to make LLMs accessible and adaptable to a wide range of mathematical inquiries, from basic problem solving to advanced theorem proving.
@@
Sachan et al. (Saxton et al., 2019) introduce a framework for evaluating neural architectures through a suite of math problems, highlighting the importance of algebraic generalization. This is complemented by GEOS (Seo et al., 2015), an innovative system that solves SAT geometry problems by integrating text understanding with diagram interpretation. The introduction of datasets and systems like Inter-GPS (Lu et al., 2021a), IconQA (Lu et al., 2021b), and PGDP5K (Hao et al., 2022; Zhang et al., 2022, 2023b) further advances the field, offering new benchmarks for abstract diagram understanding and visual language reasoning. Additionally, Scharpf et al. (Scharpf et al., 2022) indicate that unsupervised methods for formula labeling contribute to the automated understanding of mathematical documents. Collectively, these works demonstrate the evolving landscape of MQA, where the synergy between natural language processing, computer vision, and symbolic reasoning opens new avenues for educational technology and AI research.
2.2.2. Theorem Proving